OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = x - 3.

OpenStudy (anonymous):

y-4=x-6 y=x-2

OpenStudy (anonymous):

y = 1/3x + 2?

OpenStudy (anonymous):

if its parallel to y=x-3, its gradient has to be 1

OpenStudy (anonymous):

y-y1 = m(x-x1)